Use implicit defferentiation to determine the derivative of the equation of the eclipse below: 
(X/a)^2+(y/b)^2=1




1
Expert's answer
2021-06-22T14:43:15-0400
"(x\/a)^2+(y\/b)^2=1"

Differentiate both sides with respect to "x"


"\dfrac{d}{dx}((\dfrac{x}{a})^2+(\dfrac{y}{b})^2)=\dfrac{d}{dx}(1)"

Use the Chain Rule


"\dfrac{2x}{a^2}+\dfrac{2y}{b^2}\cdot\dfrac{dy}{dx}=0"

Solve for "\dfrac{dy}{dx}"

"\dfrac{dy}{dx}=-\dfrac{b^2x}{a^2y}"